The local housing supply store, Stowes, is considering adding a child care center to their store. The idea is that the child care center will attract parents to the store to do extended shopping while not having to worry about their children. The company CFO would like to test a prototype in the Athens, Georgia branch of the store. The company will test the concept for a 3-year period. Stowe’s has a marginal tax rate of 30%, and a cost of capital of 10%. The CFO believes the child care center, without considering impact on the rest of Stowe’s, will have the following financial impacts:

YEAR

0

1

2

3

Sales

$0

$0

$0

$0

Non-depreciable expenses

$275,000

$275,000

$275,000

Depreciation

$30,000

$30,000

$30,000

Investment in NWC

$10,000

Investment In Gross PPE

$90,000

At the end of year 3, the company will recover the investment in NWC and sell assets in the childcare center for a NSV of $15,000.

Stowe’s has an after-tax operating margin on sales of 40%. For the proptotype to break even, what is the yearly increase needed in sales over the three year period? (side effect)
Answer #1
Stowe's
At Break even incremental sales the
NPV of the prtotype will be zero.
Given :
Tax rate 30%
Cost of Capital 10%
NPV calculation:
PV anniuty factots @10% for 3 years 2.487
PV factor @10% for 3 yrs 0.751
Initial Investment Amt $ PV Factor/PV Annuity factor PV of cash flow
Investment in PPE                                   (90,000)                          1                            (90,000)
Investment in NWC                                   (10,000)                          1                            (10,000)
a Total Initial Investement                          (100,000)
Cash flow from operation
Assume the incremental sales for
Break even =x
Sales less non depreciation exp=x-275000
Post tax earning before depreciation=(x-275000)*0.7
Tax shield on Depreciation=30000*0.3=9000
Total Cash flow from operations= (x-275000)*0.7+9000
b Net Incremental cash flow= (x-275000)*0.7+9000                  2.487 =2.487*[(x-275000)*0.7+9000]
Terminal cash flow
Return of NWC                                    10,000                  0.751                                7,510
After tax Salvage value =15000*0.70=                                    10,500                  0.751                                7,886
c Total Terminal cash flow                              15,396
d NPV =PV of all cash flows=a+b+c=0
so , -100,000+2.487*[(x-275000)*0.7+9000]+15396=0
x=310,740.71
So incremental sales =$310,740.71
Therefore incremental Break even sales =$310,740.71 per year